OpenStudy (anonymous):

What is the sum of the first n terms of this sequence? 3,4,5,6,7...

OpenStudy (anonymous):

I'll give a medal and fan :)

OpenStudy (solomonzelman):

you know the difference right? For how many terms do you need the sum ?

OpenStudy (solomonzelman):

(last term + first term) times (number of terms / 2 ) for any arithmetic sequence.

OpenStudy (solomonzelman):

Post the options please, if you have any ????

OpenStudy (anonymous):

Sure! One sec

OpenStudy (anonymous):

B.) 3n(n+1) over 2, C. n(n+1) over 2 D. n(n+1) E. 5(n(n+1)divided by 2

OpenStudy (anonymous):

I'll try to rephrase the choices..

OpenStudy (anonymous):

k

OpenStudy (anonymous):

It might be better to write each choice.

OpenStudy (anonymous):

no. it's ok.

OpenStudy (anonymous):

whts A?

OpenStudy (anonymous):

n(n+1)/2 (the entire equation is divided by 2.)

OpenStudy (anonymous):

god. wht math u in?

OpenStudy (anonymous):

i keep getting just 25.

OpenStudy (anonymous):

I'm taking PLATO for Algebra 1. It really sucks.. I'm a senior and if I don't pass these questions, I don't graduate >.>

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

good luck;)

OpenStudy (accessdenied):

A. and C. are the same? (n (n+1)) over 2 I tried this a few times and did not get an answer from those answer choices. Could you rewrite and make sure?

OpenStudy (anonymous):

A and C are the same EXCEPT that,at the end of the equation for C,there is a '+ n'

OpenStudy (anonymous):

So it would be (n((n+1)) over 2, and 'n' is added to the entire equation.

OpenStudy (accessdenied):

You gave the sequence: 3,4,5,6,7... Your first term is 3. We can find the general kth term by induction, the common difference being 1 1st term: 3 = 3 + 0*1 2nd term: 4 = 3 + 1*1 3rd term: 5 = 3 + 2*1 kth term: = 3 + (k-1) *1 = 2 + k The last term is the nth term, so k=n; 2 + n. That means we can use our formula: S = (a_1 + a_n) * #terms / 2 = (3 + (2+n)) * n/2 But I didn't see one of the answer choices match this...

OpenStudy (anonymous):

Sorry for the confusion :/

OpenStudy (accessdenied):

No worries. Did you at least get the answer or was it not an answer choice to begin with?

OpenStudy (anonymous):

I guessed. It wasn't the right thing to do, but whatever. It was too hard phrasing the answers. Thanks for your help,though.